GATEWAY MEDIATED INTERSYSTEM CROSSING IN C_{2}H_{2}

KEVIN L. CUNNINGHAM, STEPHEN DRUCKER, and \underline{ROBERT W. FIELD}, Department of Chemistry, Massachusetts Institute of Technology, Cambridge MA 02139-4307; SUSAN J. HUMPHREY, CHRISTOPHER G. MORGAN, and ALEC M. WODTKE, Department of Chemistry, University of California at Santa Barbara, Santa Barbara CA 93106.

Laser Excited Metastable (LEM) spectra of C_{2}H_{2} indicate that a single vibrational level of the T_{3} surface couples the 3ν_{3} level of the S_{1} surface to a dense manifold of T_{1} and T_{2} states. In the proposed coupling mechanism, Gateway Mediated Intersystem Crossing (GMISC), the spin-orbit matrix element between the 3ν_{3} S_{1} state and one near-degenerate vibrational level of the T_{3} state is much larger than the direct interaction between 3ν_{3} S_{1} and any other T_{1} or T_{2} level. This is due to a near degeneracy between the S_{1} and T_{3} vibrational states and a nearby S_{1}-T_{3} surface crossing. The S_{1} state interacts with the background triplet states indirectly through T_{3} via T_{3} ∼ T_{2,1} matrix elements. Evidence for this mechanism includes a shift between the LIF and LEM spectra, an interference effect near J = 5 in the LEM spectrum, and the strength of the LEM signal for the 3ν_{3} band relative to other bands in the Ã-X̃ band system. Other experiments in which S_{1} ∼ (T_{3,2,1}, S_{0}) interactions were observed will be discussed in the context of the GMISC model.